Question 5 (0.5 points)

Use the data in the table below to answer the question.

Number of Visits

0

1,000

2,000

3,000

Total Cost

$100,000

$130,000

$160,000

$190,000

What is the average variable cost (on a per visit basis) if the number of visits is 3,000?

Question 5 options:

$ 30

$ 50

$ 63

$ 100
Answer #1

total variable cost at 3000=190000-100000=90000

What is the average variable cost (on a per visit basis) if the number of visits is 3,000=90000/3000=30 (answer)

the above will be answer..